OpenStudy (anonymous):

Find the area of the triangle with base 1 3/4 and height 4/9 yard (hint the area of a triangle equals half the product of it's base and height)

OpenStudy (zzr0ck3r):

\[(1/2)(1+3/4)(4/9) = (1/2)(7/4)(4/9) = \frac{1*7*4}{2*4*9}=?\]

OpenStudy (anonymous):

28/72

OpenStudy (zzr0ck3r):

divide top and bottom by 4

OpenStudy (anonymous):

7/18

OpenStudy (zzr0ck3r):

\[\frac{1*7*4}{2*4*9}=\frac{1*7}{2*9}\]

OpenStudy (zzr0ck3r):

make use of canceling stuff to make things easy

OpenStudy (zzr0ck3r):

correct

OpenStudy (anonymous):

\[so the answer will be 7/18yd ^{2} \]

OpenStudy (anonymous):

so the answer will be *

Directrix (directrix):

@bherre12 Assuming that the base is 1 3/4 yards and not some other unit of measure, then the answer above of 7/18 square yards is correct.